Problem 13-2A Ratios, common-size statements, and trend percents LO P1, P2, P3

[The following information applies to the questions displayed below.]

Selected comparative financial statements of Korbin Company follow:

KORBIN COMPANY
Comparative Income Statements
For Years Ended December 31, 2017, 2016, and 2015
2017 2016 2015
Sales $ 496,208 $ 380,136 $ 263,800
Cost of goods sold 298,717 237,965 168,832
Gross profit 197,491 142,171 94,968
Selling expenses 70,462 52,459 34,822
Administrative expenses 44,659 33,452 21,895
Total expenses 115,121 85,911 56,717
Income before taxes 82,370 56,260 38,251
Income taxes 15,321 11,533 7,765
Net income $ 67,049 $ 44,727 $ 30,486


KORBIN COMPANY
Comparative Balance Sheets
December 31, 2017, 2016, and 2015
2017 2016 2015
Assets
Current assets $ 52,317 $ 40,932 $ 54,716
Long-term investments 0 900 4,290
Plant assets, net 100,212 106,279 63,704
Total assets $ 152,529 $ 148,111 $ 122,710
Liabilities and Equity
Current liabilities $ 22,269 $ 22,069 $ 21,474
Common stock 67,000 67,000 49,000
Other paid-in capital 8,375 8,375 5,444
Retained earnings 54,885 50,667 46,792
Total liabilities and equity $ 152,529 $ 148,111 $ 122,710

Problem 13-2A Part 1

Required:
1. Complete the below table to calculate each year's current ratio.

2. Complete the below table to calculate income statement data in common-size percents. (Round your percentage answers to 2 decimal places.)

For each ratio listed, identify whether the change in ratio value from 2016 to 2017 is usually regarded as favorable or unfavorable.
Answer #1

1. Current Ratio = Current Assets / Current Liabilities
2017 = 52317/22269 = 2.35
2016 = 40932/22069 = 1.85
2015 = 54716/21474 = 2.55

Change in ratio from 2016 to 2017 is favorable since it increases

2.

2017 Common Size 2016 Common Size 2015 Common Size
Sales 496208 100.00% 380136 100.00% 263800 100.00%
Cost of goods sold 298717 60.20% 237965 62.60% 168832 64.00%
Gross profit 197491 39.80% 142171 37.40% 94968 36.00%
Selling expenses 70462 14.20% 52459 13.80% 34822 13.20%
Administrative expenses 44659 9.00% 33452 8.80% 21895 8.30%
Total expenses 115121 23.20% 85911 22.60% 56717 21.50%
Income before taxes 82370 16.60% 56260 14.80% 38251 14.50%
Income taxes 15321 3.09% 11533 3.03% 7765 2.94%
Net income 67049 13.51% 44727 11.77% 30486 11.56%


Common SIze is calculated using sales of year as a base.

2017 Common Size 2016 Common Size 2015 Common Size
Assets
Current assets 52317 34.30% 40932 27.64% 54716 44.59%
Long-term investments 0 0.00% 900 0.61% 4290 3.50%
Plant assets, net 100212 65.70% 106279 71.76% 63704 51.91%
Total assets 152529 100.00% 148111 100.00% 122710 100.00%
Liabilities and Equity
Current liabilities 22269 14.60% 22069 14.90% 21474 17.50%
Common stock 67000 43.93% 67000 45.24% 49000 39.93%
Other paid-in capital 8375 5.49% 8375 5.65% 5444 4.44%
Retained earnings 54885 35.98% 50667 34.21% 46792 38.13%
Total liabilities and equity 152529 100.00% 148111 100.00% 122710 100.00%

COmmon Size is calculated by taking total assets as a base